OpenStudy (anonymous):

how would I factor 3y^2 -9y -84

OpenStudy (anonymous):

\[3y^2 = 3 \times y \times y\]\[9y = 3 \times 3 \times y\]\[84 = 3 \times 7 \times 2 \times 2\]take what they have in common out and put outside brackets

OpenStudy (anonymous):

Thanks I was on the right track